# AV-Master: Dual-Path Comprehensive Perception Makes Better Audio-Visual Question Answering

### v1

@@ IV-C Ablation Studies
(a) Ablation study on main components: To explore the effectiveness of each component, we removed them individually and re-evaluated performance. As shown in Tab. VII, removing different components leads to varying degrees of performance degradation for AV-Master. Specifically, when the TDPP (Temporal Dynamic Perception Path) is removed, the performance drops to 76.83%; removing the GPAP (Global Preference Activation Path) results in a drop to 77.12%; removing the DPCL (Dual-path Contrastive Loss) lowers performance to 77.84%, and removing the AVFC (Audio-Visual Focus Capture) decreases it to 78.11%. When all components are removed simultaneously, the performance drops significantly from 78.51% to 72.94%. These results show that each component in AV-Master contributes to performance improvement, and the best results are achieved only when all components are present.
(b) Impact of template lengths: We explored the impact of the lengths of predefined templates and on model performance. As shown in Tab. IX, the average accuracy of the model reaches its highest when the template length is set to 8. If the template length is too short, important information may be missed during sampling, leading to reduced performance. Conversely, if the template length is too long, redundant information may be introduced, which can impair the model’s performance.
@@
(c) Impact of weight sharing strategies: We also explored the impact of different weight-sharing strategies in focus sampling. As shown in Tab. X, the model achieves the best overall performance when attention blocks share weights while biases remain unshared — this configuration is also adopted as our default setting. Interestingly, when only biases are shared, the model attains the lowest overall accuracy but achieves the highest accuracy on the A-QA task. This suggests that bias sharing may benefit certain subtasks while potentially hindering overall model performance.
@@
(d) Ablation study on loss functions: To analyze the contribution of each loss function, we conducted an ablation study with the results presented in the Tab. VIII. This study systematically evaluates the model’s performance by incrementally adding different loss components. The baseline model, trained only with the answer loss , achieves an average accuracy of 75.80%. Consecutively adding the contrastive loss and the preference losses and provides progressive gains. The final model, which integrates all four loss functions, achieves the highest average accuracy of 78.51%. This demonstrates that each loss component plays a vital role, and their combined effect is crucial for optimizing the model’s overall performance. It is worth noting that, as can be seen from the results in the third, fourth, and fifth rows of the table, visual preference training provides the most significant overall improvement to the model compared to other training objectives (aside from basic ). This also indicates that in the vast majority of scenarios, the model relies more heavily on visual information for question answering.
@@
(e) Impact of different feature extractors: The results in Table XI demonstrate that different visual encoders have a significant impact on the performance of AV-Master. When using ResNet-18 [59] as the visual encoder, AV-Master achieves an average accuracy of 74.04%, showing a considerable improvement compared to ST-AVQA (71.52%) with the same visual encoder. When a more powerful visual encoder, CLIP [22], is adopted, the average accuracy of AV-Master further increases to 76.31%, surpassing PSTP-Net (73.52%). Moreover, when using CLIP [22] and Internvideo2 [60] as visual encoders, AV-Master achieved the best results across all subtasks (A-QA, V-QA, AV-QA), significantly outperforming comparable methods. In summary, as the visual encoder was progressively upgraded from ResNet-18 to Internvideo2, the performance of AV-Master on all tasks steadily improved.
@@
Furthermore, the choice of audio encoders also plays a crucial role. When fixing the visual encoder to CLIP, switching from VGGish [23] to a more advanced audio encoder like CLAP [61] boosts the average accuracy from 78.51% to 79.34%, outperforming its competitor PSOT (78.42%). This trend continues with the most powerful backbones; using Internvideo2 for both visual and audio encoding, AV-Master achieves the highest overall average accuracy of 80.15%, again surpassing the competing method CoQo (79.60%). These results collectively demonstrate that the performance of AV-Master is consistently enhanced by leveraging stronger feature extractors for both the visual and audio modalities, underscoring the importance of high-quality unimodal representations for complex audio-visual reasoning tasks.
@@
(h) Ablation study on modality preference enhancement: To further investigate the effectiveness of our proposed modality preference enhancement strategy and to identify which modality has the greatest impact on overall model performance, we conduct an ablation study by selectively disabling different components during the inference stage. As shown in Table XII, removing the audio-visual preference enhancement (AVPE) leads to the most significant performance drop, reducing the average accuracy to 77.15%, which highlights the importance of jointly modeling audio-visual scenarios. Additionally, disabling visual preference enhancement (VPE) results in lower V-QA accuracy (84.02% vs. 86.58%) and overall average accuracy (77.41%), indicating that visual preference plays a critical role in AVQA. Similarly, removing the audio preference enhancement (APE) leads to degraded performance on A-QA and AV-QA, showing the necessity of modeling audio preference. Overall, the full model achieves the best performance across all metrics, confirming the complementary benefits of modality-specific preference enhancement and their joint contribution to robust AVQA performance.
(i) Impact of different enhancement methods: To explore the impact of different preference distribution fusion methods on model performance, we conduct a comparative study using three representative methods: (1) W-ADD, which first computes a weighted ratio based on the initial overall scores of the visual and auditory preference distributions before performing a weighted summation; (2) MUL, which applies element-wise multiplication between modality preferences; and (3) ADD, a simple summation of the two distributions. As shown in Table XIII, the ADD strategy achieves the best overall performance, with an average accuracy of 78.51%, outperforming both MUL (77.84%) and W-ADD (75.70%). This result suggests that a direct and unweighted summation of visual and auditory preference distributions can better preserve cross-modal complementary information and prevent overfitting to any single modality. Interestingly, although the W-ADD method introduces an adaptive weighting mechanism, its performance lags behind due to potential imbalances introduced by dynamic scaling. When viewed together with the previous ablation results in Table XII, it is evident that both the design of modality-specific preference enhancements and the fusion mechanism of these preferences play a crucial role in optimizing overall AVQA performance.
@@ IV-D Qualitative Analysis
(b) Attention visualization of different modalities: As shown in Fig. 8, we illustrate which parts of the text the visual and audio modalities attend to within the global preference activation path. For the input question “Where is the loudest instrument?”, the visual modality primarily attends to the word “instrument” while paying minimal attention to the audio-related word “loudest.” In contrast, the audio modality places significant focus on “loudest.” These results indicate that AV-Master can effectively distinguish between visual and auditory cues and accurately align them with the corresponding textual elements, which further demonstrates AV-Master’s capability in cross-modal understanding.
(c) Qualitative results comparison of different models: Fig. 9 presents a qualitative comparison among our proposed AV-Master, MLLM (VideoLLaMA2 [48] and Qwen3-Max [62]), and another AVQA expert model (QA-TIGER [6]) across four distinct audio-visual scenarios. In each case, AV-Master exhibits a superior ability to comprehend audio-visual content and provides accurate answers that align closely with both visual and auditory cues. In contrast, the other models fail to achieve such consistency. For example, in the first scenario (a), both VideoLLaMA2 and QA-TIGER incorrectly identify the instrument as an “accordion”, whereas AV-Master correctly identifies it as a “drum”, owing to its accurate interpretation of the audio signal. Similarly, in the other scenarios (c-d), AV-Master accurately infers the spatial and temporal characteristics of sound events, while the competing models fail to effectively integrate these multimodal cues. These qualitative results reinforce our quantitative findings, showing that AV-Master is more robust in fine-grained audio-visual reasoning, particularly in scenarios that require synchronized and cross-modal understanding. This further validates the effectiveness of our carefully designed modules in fully leveraging both audio and visual modality information. The inference results for VideoLLaMA2 and Qwen3-Max were obtained from its official demo, simulating application scenarios in low-resource environments.

### v2

@@ IV-C Ablation Studies
(a) Ablation study on main components: To explore the effectiveness of each component, we removed them individually and re-evaluated the performance. As shown in Tab. VIII, removing different components leads to varying degrees of performance degradation for AV-Master. Specifically, removing the TDPP (Temporal Dynamic Perception Path) causes the largest drop in overall accuracy, from 78.51% to 76.83%, indicating that fine-grained temporal perception is the most critical source of performance gain in our framework. Removing the GPAP (Global Preference Activation Path) also results in a clear decrease to 77.12%, which verifies the importance of modeling modality preference from a global perspective. In addition, removing the DPCL (Dual-path Contrastive Loss) lowers the performance to 77.84%, showing that collaborative optimization between the two paths is beneficial for learning more discriminative cross-modal representations.
We further analyze the contributions of the two modules inside the temporal dynamic perception path. Removing AVFC (Audio-Visual Focus Capture) reduces the average accuracy to 78.11%, while removing AVKF (Audio-Visual Key Fusion) yields 78.19%. These results suggest that AVFC and AVKF are complementary: the former is responsible for extracting fine-grained question-relevant cues from the original audio-visual sequence, whereas the latter further integrates the focused audio-visual features with multimodal context to produce a more discriminative fused representation. When all designed components are removed simultaneously, the performance drops substantially from 78.51% to 72.94%, confirming that the overall improvement comes from the joint contribution of all proposed modules rather than any single design alone.
Moreover, to further compare our modality modeling strategy with a representative late-fusion approach, we introduce a Dynamic Fus variant implemented based on a question-oriented audio-visual adaptive fusion strategy. This variant achieves an average accuracy of 77.41%, which is 1.10% lower than AV-Master. The performance gap is particularly evident on the A-QA and V-QA sub-tasks, where AV-Master improves from 77.53% to 79.95% and from 84.97% to 86.58%, respectively. These results indicate that assigning dynamic modality weights only at the fusion stage is less effective than our preference enhancement strategy. Our approach directly models modality preference from the initial unimodal features. By activating question-relevant global information before multimodal entanglement, GPAP provides more reliable modality guidance, which complements the fine-grained temporal cues captured by TDPP.
(b) Impact of template lengths: We further investigate the impact of predefined audio and visual template lengths on model performance by independently varying the lengths of and . As shown in Table X, symmetric and moderate-length template configurations are more favorable. For example, the model achieves the highest average accuracy when both template lengths are set to 8. A possible reason is that our AVFC module adopts a parameter-sharing strategy across modalities. Using balanced template lengths helps maintain similar representational capacity in the audio and visual branches, making the shared focus-sampling process more stable and effective. In contrast, overly short templates may miss useful cues, while overly long or highly imbalanced templates may introduce redundant information and weaken cross-modal consistency, leading to degraded performance.
@@
(c) Impact of weight sharing strategies: We also explored the impact of different weight-sharing strategies in focus sampling. As shown in Table XI, the model achieves the best overall performance when attention blocks share weights while biases remain unshared — this configuration is also adopted as our default setting. Interestingly, when only biases are shared, the model attains the lowest overall accuracy but achieves the highest accuracy on the A-QA task. This suggests that bias sharing may benefit certain subtasks while potentially hindering overall model performance.
@@
(d) Ablation study on loss functions: To analyze the contribution of each loss function, we conducted an ablation study with the results presented in the Tab. IX. This study systematically evaluates the model’s performance by incrementally adding different loss components. The baseline model, trained only with the answer loss , achieves an average accuracy of 75.80%. Consecutively adding the contrastive loss and the preference losses and provides progressive gains. The final model, which integrates all four loss functions, achieves the highest average accuracy of 78.51%. This demonstrates that each loss component plays a vital role, and their combined effect is crucial for optimizing the model’s overall performance. It is worth noting that, as can be seen from the results in the third, fourth, and fifth rows of the table, visual preference training provides the most significant overall improvement to the model compared to other training objectives (aside from basic ). This also indicates that in the vast majority of scenarios, the model relies more heavily on visual information for question answering.
@@
(e) Impact of different feature extractors: The results in Table XII demonstrate that different visual encoders have a significant impact on the performance of AV-Master. When using ResNet-18 [16] as the visual encoder, AV-Master achieves an average accuracy of 74.04%, showing a considerable improvement compared to ST-AVQA (71.52%) with the same visual encoder. When a more powerful visual encoder, CLIP [39], is adopted, the average accuracy of AV-Master further increases to 76.31%, surpassing PSTP-Net (73.52%). Moreover, when using CLIP [39] and Internvideo2 [47] as visual encoders, AV-Master achieved the best results across all subtasks (A-QA, V-QA, AV-QA), significantly outperforming comparable methods. In summary, as the visual encoder was progressively upgraded from ResNet-18 to Internvideo2, the performance of AV-Master on all tasks steadily improved.
@@
Furthermore, the choice of audio encoders also plays a crucial role. When fixing the visual encoder to CLIP, switching from VGGish [17] to a more advanced audio encoder like CLAP [49] boosts the average accuracy from 78.51% to 79.34%, outperforming its competitor PSOT (78.42%). This trend continues with the most powerful backbones. When using Internvideo2 for both visual and audio encoding, AV-Master achieves the highest overall average accuracy of 80.15%, again surpassing the competing method CoQo (79.60%). These results collectively demonstrate that the performance of AV-Master is consistently enhanced by leveraging stronger feature extractors for both the visual and audio modalities, underscoring the importance of high-quality unimodal representations for complex audio-visual reasoning tasks.
@@
(h) Ablation study on modality preference enhancement: To further investigate the effectiveness of our proposed modality preference enhancement strategy and to identify which preference component contributes most to the final prediction, we conduct an ablation study by selectively disabling different enhancement settings during inference. APE, VPE, and AVPE denote Audio Preference Enhancement, Visual Preference Enhancement, and Audio-Visual Preference Enhancement, respectively. Specifically, APE and VPE refer to enhancing the final answer prediction by applying the audio and visual preference distributions during inference, respectively, while AVPE denotes their joint audio-visual preference enhancement effect. As shown in Table XIII, removing the audio-visual preference enhancement (AVPE) leads to the most significant performance drop, reducing the average accuracy to 77.15%, which highlights the importance of jointly modeling audio-visual scenarios. Additionally, disabling visual preference enhancement (VPE) results in lower V-QA accuracy (84.02% vs. 86.58%) and overall average accuracy (77.41%), indicating that visual preference plays a critical role in AVQA. Similarly, removing the audio preference enhancement (APE) leads to degraded performance on A-QA and AV-QA, showing the necessity of modeling audio preference. Overall, the full model achieves the best performance across all metrics, confirming the complementary benefits of modality-specific preference enhancement and their joint contribution to robust AVQA performance.
(i) Impact of different enhancement methods: To explore the impact of different preference distribution fusion methods on model performance, we conduct a comparative study using three representative methods: (1) W-ADD, which first computes a weighted ratio based on the initial overall scores of the visual and auditory preference distributions before performing a weighted summation. (2) MUL, which applies element-wise multiplication between modality preferences. (3) ADD, a simple summation of the two distributions. As shown in Table XIV, the ADD strategy achieves the best overall performance, with an average accuracy of 78.51%, outperforming both MUL (77.84%) and W-ADD (75.70%). This result suggests that a direct and unweighted summation of visual and auditory preference distributions can better preserve cross-modal complementary information and prevent overfitting to any single modality. Interestingly, although the W-ADD method introduces an adaptive weighting mechanism, its performance lags behind due to potential imbalances introduced by dynamic scaling. When viewed together with the previous ablation results in Table XIII, it is evident that both the design of modality-specific preference enhancements and the fusion mechanism of these preferences play a crucial role in optimizing overall AVQA performance.
@@ IV-D Qualitative Analysis
(b) t-SNE comparison between preference enhancement and dynamic fusion: As shown in Fig. 7, we provide a t-SNE visualization on the MUSIC-AVQA dataset to compare the learned representations of the proposed audio-visual preference enhancement strategy and the Dynamic Fus variant. It can be observed that the features produced by our preference enhancement strategy exhibit more compact intra-class clusters and clearer inter-class boundaries. In particular, several instrument categories show more concentrated distributions with less overlap with neighboring classes, indicating stronger discriminative capability of the learned representations. In contrast, the Dynamic Fus variant presents relatively more scattered feature distributions and more pronounced overlap among some categories, suggesting that estimating modality importance only after multimodal fusion is more susceptible to interference from entangled cross-modal information. This can be attributed to the fact that our method, under question guidance, directly models modality preference from the initial audio and visual features, rather than inferring modality weights only at the fusion stage. Such a design activates question-relevant global information while preserving modality-specific semantics, enabling the model to learn a more structured and separable feature space.
(c) Attention visualization of different modalities: As shown in Fig. 9, we illustrate which parts of the text the visual and audio modalities attend to within the global preference activation path. For the input question “Where is the loudest instrument?”, the visual modality primarily attends to the word “instrument” while paying minimal attention to the audio-related word “loudest.” In contrast, the audio modality places significant focus on “loudest.” These results indicate that AV-Master can effectively distinguish between visual and auditory cues and accurately align them with the corresponding textual elements, which further demonstrates AV-Master’s capability in cross-modal understanding.
(d) Qualitative results comparison of different models: Fig. 10 presents a qualitative comparison among our proposed AV-Master, MLLM (VideoLLaMA2 [6] and Qwen3-Omni [51]), and another AVQA expert model (QA-TIGER [19]) across four distinct audio-visual scenarios. In each case, AV-Master exhibits a superior ability to comprehend audio-visual content and provides accurate answers that align closely with both visual and auditory cues. In contrast, the other models fail to achieve such consistency. For example, in the first scenario (a), both VideoLLaMA2 and QA-TIGER incorrectly identify the instrument as an “accordion”, whereas AV-Master correctly identifies it as a “drum”, owing to its accurate interpretation of the audio signal. Similarly, in the other scenarios (c-d), AV-Master accurately infers the spatial and temporal characteristics of sound events, while the competing models fail to effectively integrate these multimodal cues. These qualitative results reinforce our quantitative findings, showing that AV-Master is more robust in fine-grained audio-visual reasoning, particularly in scenarios that require synchronized and cross-modal understanding. This further validates the effectiveness of our carefully designed modules in fully leveraging both audio and visual modality information. The inference results for VideoLLaMA2 and Qwen3-Omni were obtained from its official demo, simulating application scenarios in low-resource environments.